Electronic Transactions on Artificial Intelligence^{Vol. 5 (2001), Section C: pp. 15-63}^{http://www.ep.liu.se/ej/etai/2001/016/}^{http://www.ep.liu.se/ej/etai/2001/016/}  Refereed Article

J. Hefflin, J. Hendler, and S. Luke: SHOE: A Prototype Language for the Semantic Web

Author’s affiliation: Abstract:

Publication and review history

This work was first presented at the Workshop on Semantics for the Web, held at Schloss Dagstuhl. Attendance at this workshop was based on invitations. Speakers at the workshop were invited to submit their articles for consideration in the forthcoming workshop volume. The subsequent publication and review history for the present article was as follows.

(To be added)

The review policy and the quality requirements for acceptance are documented at http://www.etaij.org/info/

Article maintenance

The Electronic Transactions on Artificial Intelligence maintains an Article Interaction Page for this article at

http://www.etaij.org/ra/seweb/004/ Besides the publication and review history, it contains links for contacting the author(s), as well as to amendments and other related information for the article. It is intended to keep these links up-to-date in the future.

Copyright conditions

For all versions of the article mentioned above, the copyright belongs to the author. The publishing agreements specify that it is permitted for anyone to download the article from the net, to print out single copies of it, and to use classroom sets of copies for academic purposes. Please refer to the article URL:s for additional conditions.